The sides of a quadrilateral are all positive integers and three of them are 10,15,20. How many possible values are there for fourth side?

Grade:10

1 Answers

Saurabh Koranglekar
4 years ago
Property of quadrilateralis all sides of quadrilaterial is less than sum of all other sides
sides are of length
5, 10, 20
so maximum length can be, 5+10+20-1= 34 ( as it's less than not equal to sum of all other sides)
and minimum can be, let it be x
20=10+5 +x
=>x=5
x=6 ( as it's less than not equal to sum of all other sides)
so, from 6 to 34 interger can be
34-5=29answer